OpenStudy (anonymous):

Ill fan those who can help me with these and Ill give medal to the best answer. Find the area of an equilateral triangle whose apothem is square root of 48.

OpenStudy (anonymous):

@goformit100 @StudyGurl14 @Starbucks @Squirrels @Skyz @ShadowLegendX

OpenStudy (anonymous):

@PRAETORIAN.10 @prowrestler @peekaboopork @Perce @powerstroke97 @pondamo

OpenStudy (studygurl14):

Maybe this will help. I really don't know about the relationship between apothems and the area of triangles. http://www.vitutor.com/geometry/plane/equilateral_triangle.html

TheSmartOne (thesmartone):

Do you have a picture?

OpenStudy (anonymous):

None. The problem says it all

OpenStudy (confluxepic):

Ok.

OpenStudy (lyrae):



OpenStudy (anonymous):

I think it is is 1/2(4rad3)(1),

OpenStudy (anonymous):

a=1/2ap apothom simplified 4rad3, perimeter 1, and total answer 3.5 in decimal form

OpenStudy (anonymous):

how did you get that @n648c788

OpenStudy (anonymous):

@Lyrae help me please :(

OpenStudy (anonymous):

Will you elaborate it @n648c788

OpenStudy (anonymous):

formula A=1/2ap a=apothom p=perimeter so i simply rad48 to 4rad3 and since perimeter is 1 A=1/2(4rad3)(1)

OpenStudy (anonymous):

If I simply all way to decimal form 3.5.

OpenStudy (anonymous):

it would be8rad3 if i just times 1/2 and left it in radical

OpenStudy (anonymous):

Is it making sense?

OpenStudy (anonymous):

why would you use rad?

OpenStudy (anonymous):

SO the answer is 3.5

OpenStudy (anonymous):

it depends what form your teacher specify the answer to be in

OpenStudy (anonymous):

yes in decimal form

OpenStudy (anonymous):

he's just seeking for the area . I dont get it why you use rad..? RADIAN?

TheSmartOne (thesmartone):

rad is the same thing as square root

TheSmartOne (thesmartone):

rad is radical sign...

OpenStudy (anonymous):

no radical for like square root 3

OpenStudy (anonymous):

radians is measure of ange like pi/2

OpenStudy (anonymous):

just notify me if you need any more help

OpenStudy (anonymous):

my calculator says its 1.31 thingy is there something wrong with my input?"

OpenStudy (anonymous):

type 1/2(4rad3) in calculator

TheSmartOne (thesmartone):

\[\frac{ 1 }{2}\times4\sqrt{3}\] this is what @n648c788 is sayng...

OpenStudy (anonymous):

1/2(4sqrt3)

OpenStudy (anonymous):

yes

OpenStudy (lyrae):

How did you get perimeter 1 with apthem 4 rad 3? One side of the triangle should be larger than the apotem, since the three sides have equal length shouldent the perimeter be grater than atleast tree times the apothem?

OpenStudy (anonymous):

everything ok?

TheSmartOne (thesmartone):

I wasn't sure about the perimeter being 1 either @Lyrae

OpenStudy (anonymous):

Yaaaa @n648c788 atleast you help me not the other one i just met earlier who promised to help me lol..

TheSmartOne (thesmartone):

Becasue it would be a 30-60-90 triangle you would figure out half of one side and then double it to find the whole side of the triangle...

OpenStudy (anonymous):

since it was not given you can assume it to be 1

OpenStudy (anonymous):

Time to wrap this up. I have 4 more questions I guess.. to ask xD